A demonstration gyroscope wheel is constructed by
Answer #1

The vertical forces must sum to zero.

Denote the upward forces that the hands exert as FL and FR. \small \tau = (FL − FR )r, where
r = 0.210 m.

The conditions that FL and FR must satisfy are \small F_{L}+F_{R}=w and \small F_{L}-F_{R}=\Omega \frac{I\omega }{r} , where the
second equation is \small \tau =\Omega L , divided by r. These two equations can be solved for the forces by first adding
and then subtracting, yielding \small F_{L}=\frac{1}{2}\left ( w+\Omega \frac{I\omega }{r} \right ) and \small F_{R}=\frac{1}{2}\left ( w-\Omega \frac{I\omega }{r} \right ) . Using the values

w=mg = (6 kg)(9.8 m/s2) = 58.8 N and

Radius of the bicycle wheel, a = 0.325 m

\small \frac{I\omega }{r}=\frac{6 \times 0.325^{2} \times 5.30 \times 2\pi}{0.210}=100.5 kg.m/s

Gives

FL = 29.4 N + \small \Omega (50.25N.s), FR = 29.4 N - \small \Omega (50.25N.s).

Part C

To find the force each hand exerts on the shaft when the shaft is rotating in a horizontal plane about its center at 0.300 rev/s

\small \Omega = 0.3 rev/s =1.89 rad/s, FL = 124.37 N, FR = − 65.57 N, with the minus sign indicating a downward
force.

Part D

To find the rate of the shaft rotates in order that it may be supported at one end only

FR = 0, gives  \small \Omega =\frac{29.4}{50.25}=0.585 rad/s, which is 0.0931 rev/s.

\small \Omega =0.0931 rev/s